# EZIGen: Enhancing zero-shot personalized image generation with precise subject encoding and decoupled guidance

### v3

@@ Abstract
Zero-shot personalized image generation models aim to produce images that align with both a given text prompt and subject image, requiring the model to effectively incorporate both sources of guidance. However, existing methods often struggle to capture fine-grained subject details and frequently prioritize one form of guidance over the other, resulting in suboptimal subject encoding and an imbalance in the generated images. In this study, we uncover key insights into achieving high-quality balances on subject identity preservation and text-following, notably that 1) the design of the subject image encoder critically influences subject identity preservation, and 2) the text and subject guidance should take effect at different denoising stages. Building on these insights, we introduce a new approach, EZIGen, that employs two main components: a carefully crafted subject image encoder based on the pre-trained UNet of the Stable Diffusion model, following a process that balances the two guidances by separating their dominance stage and revisiting certain time steps to bootstrap subject transfer quality. Through these two components, EZIGen achieves state-of-the-art results on multiple personalized generation benchmarks with a unified model and 100 times less training data. Demo Page: Demo.
@@ 1 Introduction
Personalized image generation methods enable users to create images by combining text prompts with subject images, following the principle of ‘my subject’ following ‘my instructions’. Existing solutions fall into two categories: test-time tuning-based [11, 32, 1, 19, 20, 13] and zero-shot inference-based[39, 28, 30, 7]. Test-time tuning involves fine-tuning model parameters and introducing subject tokens, allowing detailed control but requiring time-consuming re-training. In contrast, zero-shot methods generate images directly from the given subjects without re-training, offering greater efficiency. This paper focuses on improving zero-shot methods.
Existing zero-shot methods [39, 7, 30, 28] predominantly focus on the transfer of a reference subject’s appearance into the generated image. These methods typically first encode the subject image and then integrate the encoded features into the UNet of the diffusion model. There are several existing options for subject image encoders. ELITE [39] and Subject Diffusion [28] utilize CLIP image encoders, while AnyDoor [7] adopts DINOv2 to achieve better feature map extraction. However, these methods all fail to capture details of the subject image, leading to significant distortion compared to the subject image. Recently, BootPIG [30] employs a Reference UNet model [15] as a subject feature extractor for common objects, reasoning that the feature space of a UNet would be more aligned with the feature space used in image generation. Although this approach is reasonable, it leaves many design aspects unexplored, such as which timestep to use and detailed extractor model configuration. Following their steps, our research reveals that those aspects might have a big impact on the identity preservation capability, as shown in Fig. 2.
Another often overlooked challenge in existing methods is balancing the parallel user inputs: subject image and text prompt. Although these guidance appear to be independent—the subject image focuses on preserving identity and the text prompt directs the model to follow user instructions—they in fact frequently interfere with one another. For instance, as shown in Fig. 3, both the text prompt “a dog in police outfit” and the subject dog image define the dog’s appearance to some extent simultaneously. In such cases, prior works either prioritize identity preservation at the expense of text coherence [28], successfully follow the text but struggle to maintain subject identity [30], or perform suboptimally in both aspects [39, 20].
@@
In this paper, we address the aforementioned challenges and Enhance Zero-shot personalized Image Generation by proposing a novel method called EZIGen. Firstly for subject encoding, we follow [30] to initialize a Stable Diffusion UNet as an extractor to achieve good feature alignment between the subject and generated images. However, our unique contribution lies in identifying the ‘devils in the details’: Different from [15, 6] that removes the time embedding and fully-finetunes a Stable Diffusion UNet to serve as time-irrelevant Reference UNets and [30] where the Reference UNet is tuned and is synchronized with the denoising timesteps, we discover that simply adding noise and denoising the subject image to the final denoising step $T_{1\rightarrow 0}$ with a vanilla frozen Stable Diffusion UNet yields significantly improved subject representations. Secondly, to better balance subject identity and text adherence, we decouple the generation process into two distinct stages: the Sketch Generation Process, which forms a coarse sketch latent from text prompts, and the Appearance Transfer Process, which injects the encoded subject details via the adapter. This decoupling explicitly separates guidance signals. Additionally, we observe that the sketch latent can impact the quality of appearance transfer—a sketch latent closer to the subject tends to produce better results. Therefore, we introduce an iterative pipeline that repeatedly converts the generated image back into the editable noisy latent to serve as a new sketch latent that ensembles more visual cues from the given subject, and progressively achieves satisfactory results as the iteration continues. With the aforementioned designs, our contribution can be summarized as follows: 1) We identify that the design choices of the subject image encoder significantly impact identity preservation and propose an improved design over existing solutions. 2) We propose a solution that combines guidance decoupling with iterative bootstrapping to address the guidance balance issue. 3) We extend our approach to personalized image editing and domain-specific personalization tasks. 4) Extensive experiments and analysis demonstrate the state-of-the-art performance of EZIGen.
@@ 2 Related Works
2.2 Tunning-based personalized image generation.
@@ 2.3 Zero-shot personalized image generation
To tackle the aforementioned issues, some researchers introduced zero-shot methods that accept new subjects without requiring retraining. Some attempts [33, 40, 23, 42, 38], developed zero-shot personalized generation techniques for domain-specific data, such as human faces or other constrained applications. Recently, for more general objects, ELITE[39] pioneeringly projects the image into text space and refines with detailed patch features, BLIP-Diffusion[20] leverages a Q-Former[21] to align text and images, Subject-Diffusion[28] inject grounding info to control the subject position, BootPIG[30] setups a Reference UNet to extract subject features. Nevertheless, these methods struggle with issues like degraded subject-identity preservation due to sub-optimal feature extractor utilization or poor balancing between the text and subject guidance during the generation process.
@@ 3.1 Encoding and injecting subject feature
The identity information of the subject image is often extracted using an image encoder. Current methods use options like CLIP[28], DINO-V2[7]), or a Reference UNet[30] trained based on Stable Diffusion UNet. Leveraging a Stable Diffusion UNet for subject feature extraction is particularly appealing as it naturally aligns with the Stable Diffusion feature space, reducing the challenge of aligning disparate feature spaces. However, this design raises open questions about configuring the feature extractor and integrating subject encoding into the Stable Diffusion UNet. In BootPIG [30], all Reference UNet parameters are fine-tuned, with progressively noised input images fed at each timestep, allowing subject image features to be injected into its Main UNet for generation. However, we identify potential issues: overly-noised features may distort subject information, and tuning Stable Diffusion parameters into a Reference UNet can lead to suboptimal subject encoding, leading to degraded subject encoding(Fig. 2 shows some examples). In this paper, we propose a simple yet effective solution: First, we directly take the vanilla Stable Diffusion UNet and conduct one-step denoising on a one-step noised subject image to obtain intermediate subject feature maps from UNet as the subject appearance feature. Second, we again take the same UNet and insert cross-attention adapters in its transformer block to adopt the subject image guidance.
@@ 3.3 Iterative Appearance Transfer
Based on the above analysis and empirical results, we observe that when the sketch latent $x_{t}$ and the subject share similar semantics, such as rough color patterns, the appearance transfer process improves as the adapter establishes a stronger matching within the UNet’s noisy latent space and could thus transfer more subject appearances. To enhance this, we introduce an iterative generation scheme: when we obtain the clean latent $x_{0}$ after the Appearance Transfer Process, we add noise with level $t$ to $x_{0}$ to obtain a new $x_{t}$ as the editable sketch latent for another round of the Appearance Transfer Process, as shown in the bottom part of Fig. 4. This process repeats until the similarity between the newly generated image and the previous image exceeds a predefined threshold, indicating that the appearance transfer is complete and no further information is added. Integrating the designs above, our model successfully balances subject identity preservation and text adherence, ensuring comprehensive guidance-adhering without compromise.
@@ 4.1 Implementation details
Benchmark and evaluation. We mainly evaluate EZIGen on two standard benchmarks: DreamBench [32] for personalized image generation, DreamEdit [22] for personalized image editing. For both tasks, we follow Subject Diffusion’s protocol, averaging scores over 6 random runs. Regarding the evaluation metric, we report CLIP-T and DINO scores to assess text adherence and subject identity on DreamBench and we evaluate DINO-sub for foreground subject similarities on DreamEditBench, using SAM [18] for mask extraction. We abandon the traditional CLIP-I scores and regard the DINO score as the main metric for subject similarity, as they better depict the detailed patch-level differences between images and are more sensitive to subject quality degradation, as discussed in [32, 38], please refer to Appendix C for details. Besides standard evaluation, we additionally validate EZIGen on the challenging human content generation benchmark [40] to test the ability to generalize to domain-specific generation tasks, which usually require special design or in-domain training while EZIGen does not. We report ID preservation and prompt consistency following [40, 28]. At last, we involve human evaluation for the comparisons mentioned above, the percentage scores denote the frequency of the methods’ generation results being chosen as the best within the given sample batch under specific criteria.
Training dataset construction. We create training pairs from COCO2014 and YoutubeVIS. In COCO2014, 1-4 objects are cropped from the target image as subject images, while in YoutubeVIS, we extract images of the same subject from different frames as pairs, following [7]. We sampled 100k pairs from each dataset, totaling 200k pairs. Please refer to Fig. 13 for more details and dataset visualization.
Experiment settings. For a fair comparison, we utilize SD2.1-base as the base model, with an image resolution of 512×512. The adapter is initialized from the self-attention module within each transformer block to enhance compatibility. We train the model for 1 epoch using a batch size of 1, with the Adam optimizer and a learning rate of $51e-$ 5. During inference, iterations are designed to stop automatically when the newly generated image exhibits a sufficiently high similarity with the image from the previous loop, ensuring efficient convergence and maintaining generation quality. Please refer to Appendix B for more details.
@@
  Flex. ID Overall Quality
3.1% 2.2% 1.1%
15.4% 17.3% 22.2%
6.9% 7.4% 9.0%
12.3% 2.7% 2.3%
11.2% 9.1% 6.9%
20.7% 15.1% 18.4%
30.4% 46.2% 40.1%
@@ 4.2 Method comparisons on DreamBench dataset
In Tab. 1 and Fig. 5, we compare our method with existing zero-shot personalized image generation approaches on the DreamBench dataset [32], including tuning-based methods for reference. We also present the number of reference images required for each subject and the training dataset size to highlight the cost of each method. The results show that our base model achieves state-of-the-art performance text-following, subject identity preservation, and balance among all methods. Compared to the previous state-of-the-art, Subject Diffusion [28], our model outperforms in both CLIP-T (0.316 vs. 0.293) and DINO score (0.718 vs. 0.711), demonstrating superior flexibility in generated images and higher subject identity preservation, without sacrificing balance. Moreover, this was accomplished with 100$\times$ less training data, which we attribute to our simplified design and the use of the original Stable Diffusion UNet. These results highlight the advantages of using a Stable Diffusion UNet over CLIP for feature extraction, alongside our decoupling design that improves text coherence. When compared with BootPIG [30], our method surpasses it in text-following ability (0.316 vs. 0.311) and significantly outperforms in DINO score (0.718 vs. 0.674), while requiring fewer subject images. This is due to our effective utilization of the Stable Diffusion UNet extractor. Additionally, our method outperforms previous open-source zero-shot methods [39, 20] across all evaluation metrics.
@@ 4.3 Validation on personalized image editing task
As outlined in Sec. 3.4, our method adapts well to personalized image editing by replacing noise addition with image inversion [27] in the Appearance Transfer Process. We compare our model against previous state-of-the-art methods on the DreamEdit benchmark, as shown in Tab. 2 and Fig. 6. Our method significantly outperforms others, primarily due to our advanced subject feature utilization, where the original Stable Diffusion UNet produces high-quality, detailed intermediate appearance features, and the adapter accurately matches them to the sketch latent. Notably, even though MimicBrush [6] leverages a tuned Reference UNet and an extra CLIP image encoder for fine-grained feature extraction, it sometimes struggles to modify areas with large semantic differences from their surroundings. For example, in the last column of Fig. 6, the model has difficulty adjusting the dog face in contrast to the astronaut suit. We attribute this limitation to their masked image modeling training strategy, which aims to recover masked areas based on neighboring tokens.
@@ 4.4 Ablation study
Effectiveness of our subject encoder. To evaluate the quality of the encoded subject feature, we start with a fully noised sketch latent (i.e. random noise) and apply subject guidance at all timesteps, dubbed as “full injection”. As shown in Fig. 8 experiments 1) and 2), compared to the original text-guided generation, the encoded subject representation resembles the subject effectively, achieving a DINO score of 0.762. However, the strong subject features override the weaker sketch latent in early timesteps, leading to severe copy-paste effects and a low CLIP-T score of 0.286.
Figure 7: Qualitative result of the ablation study.
@@
Figure 9: Visualization of subject ID preservation performance under various UNet configurations.
  #ID Train SDU Denoise Timestep DINO
1) BootPIG* Adaptive 0.589
2) Adaptive 0.615
3) $t^{\text{sub}}$=1 0.668
4) $t^{\text{sub}}$=800 0.427
5) $t^{\text{sub}}$=500 0.598
6) $t^{\text{sub}}$=300 0.661
7) $t^{\text{sub}}$=100 0.725
8) Ours $t^{\text{sub}}$=1 0.762
Figure 10: Scores evaluating subject ID preservation under different SDU configurations. The best result is obtained by using the fixed SDU and using a small denoising timestep $t^{\text{sub}}$. ”BootPIG*” indicates BootPIG’s training strategy.
4.5 Comparing feature extractor with other choices
We evaluate the effectiveness and efficiency of our subject extraction compared to CLIP, DINO, and Reference UNet based methods — Subject Diffusion [28], AnyDoor [7], and BootPIG [30]—as the current best practices for each feature extraction paradigm respectively. For fair comparisons, we ablate the models for AnyDoor and Subject Diffusion (lines 2 and 4) so that subject information is derived solely from the feature extractor. We denote Stable Diffusion UNet, our base extractor, as SDU for simplicity.
@@ 4.5 Comparing feature extractor with other choices
  Subject Encoder Training Strategies DINO Dataset Size
@@
Figure 12: Comparing subject encoders, our method is the most efficacious, achieving the highest quality with minimal data requirements. SDU denotes Stable Diffusion UNet.
Evaluating alternative feature extractor configurations. We examined the impact of tuning SDU parameters and adding varying levels of noise to subject images. Fig. 10 lists the experiment settings and their performances, where check marks/cross marks indicate whether the SDU extractor is trained or frozen, the corresponding visualization results are shown in Fig. 10. “Adaptive” means that the tuned SDU extractor shares the same denoising timestep as the main SDU, while arguments starting with “$t^{\text{sub}}$” indicate the level of noise we add to the subject image. From lines (1) and (2), we find that tuning the SDU disrupts the original generative capabilities of the SDU, resulting in a performance drop compared to freezing it. Lines (1) and (3) show that fixing the denoising timestep of the SDU to a small $t^{\text{sub}}$ produces better subject identity with clearer feature maps. By fixing the SDU and adjusting $t^{\text{sub}}$, we validated its importance; as shown in lines (4) to (8), subject identity scores increase as $t^{\text{sub}}$ decreases. This aligns with the intuition that the SDU encodes clearer features in the final timesteps of the generation process.

### v4

@@ Abstract
Zero-shot personalized image generation models aim to produce images that align with both a given text prompt and subject image, requiring the model to incorporate both sources of guidance. Existing methods often struggle to capture fine-grained subject details and frequently prioritize one form of guidance over the other, resulting in suboptimal subject encoding and imbalanced generation. In this study, we uncover key insights into overcoming such drawbacks, notably that 1) the choice of the subject image encoder critically influences subject identity preservation and training efficiency, and 2) the text and subject guidance should take effect at different denoising stages. Building on these insights, we introduce a new approach, EZIGen, that employs two main components: leveraging a fixed pre-trained Diffusion UNet itself as subject encoder, following a process that balances the two guidances by separating their dominance stage and revisiting certain time steps to bootstrap subject transfer quality. Through these two components, EZIGen, initially built upon SD2.1-base, achieved state-of-the-art performances on multiple personalized generation benchmarks with a unified model, while using 100 times less training data. Moreover, by further migrating our design to SDXL, EZIGen is proven to be a versatile model-agnostic solution for personalized generation.
@@ 1 Introduction
Personalized image generation methods enable users to create images by combining text prompts with subject images, following the principle of ‘my subject’ following ‘my instructions’. Existing solutions fall into two categories: test-time tuning-based [13, 35, 2, 21, 23, 15] and zero-shot inference-based [44, 31, 33, 8, 42, 48, 9, 46, 40, 7]. Test-time tuning involves fine-tuning model parameters for each input subject, in contrast, zero-shot methods generate images directly from the given subjects without re-training, offering greater efficiency. Among these methods, some recent attempts [46, 40] rely on heavy transformer-based models, producing satisfying results at the cost of unaffordable computational resources, reducing practicality. Thus, this study focuses on improving UNet-based zero-shot methods.
Existing UNet-based zero-shot methods [44, 8, 33, 31, 48, 43] typically rely on an image encoder to capture the subject’s appearance, and then transfer it into the target image. Some adopt a CLIP Image encoder [44, 31, 42, 48], while AnyDoor [8] adopts DINOv2 to achieve better feature map extraction. However, these methods all fail to capture details of the subject image, leading to significant distortion compared to the subject image. BootPIG recently [33] tunes a Reference UNet model [17] as a subject feature extractor, reasoning that the feature space of a UNet would be more aligned with the feature space used in image generation. Though reasonable, it leaves many aspects unexplored, such as which timestep to use and many other detailed configurations. Following their steps, our research reveals that those aspects might have a big impact on the identity preservation capability, as shown in Fig. 2.
Another often overlooked challenge in existing methods is achieving a coherent balance between the two parallel user inputs: the subject image and the text prompt. Although these inputs appear orthogonal—where the subject image is intended to preserve identity and the text prompt guides semantic composition—they often exert conflicting constraints on the generation process. As illustrated in Fig. 3, when the textual prompt specifies “a robot horse in the neon city” and the subject image contains a horse facing a different direction or with mismatched global attributes, baseline models such as [43] fail to reconcile both modalities. This results in either poor identity preservation or text-following depending on the relative dominance of either input. Notably, as the conditioning strength of the subject image ($\partial$) increases, the generated output increasingly conforms to its pose and structure but diverges from the prompt context. Conversely, low $\partial$ values lead to poor identity retention. This trade-off highlights the inherent tension in current conditioning strategies, underscoring the need for a more principled mechanism to disentangle and balance subject fidelity with semantic controllability. In such cases, prior works either prioritize identity preservation at the expense of text coherence [31], successfully follow the text but struggle to maintain subject identity [33, 43, 48, 46, 40], or perform suboptimally in both aspects [44, 23].
In this paper, we address the aforementioned challenges and Enhance Zero-shot personalized Image Generation by proposing a novel method called EZIGen. Firstly, for subject encoding, we follow [33] to initialize a Stable Diffusion UNet as an extractor to achieve good feature alignment between the subject and generated images. However, our unique contribution lies in identifying the ‘devils in the details’: Different from [17, 7] that removes the time embedding and fully-finetunes a Stable Diffusion UNet to serve as time-irrelevant Reference UNets and [33] where the Reference UNet is tuned and is synchronized with the denoising timesteps, we discover that simply adding noise and denoising the subject image to the final denoising step $T_{1\rightarrow 0}$ with a vanilla frozen Stable Diffusion UNet yields significantly improved subject representations. Secondly, to better balance subject identity and text adherence, we decouple the generation process into two distinct stages: the Sketch Generation Process, which forms a coarse sketch latent from text prompts, and the Appearance Transfer Process, which injects the encoded subject details via the adapter. This decoupling explicitly separates guidance signals. Additionally, we observe that the sketch latent can impact the quality of appearance transfer—a sketch latent closer to the subject tends to produce better results. Therefore, we introduce an iterative pipeline that repeatedly converts the generated image back into the editable noisy latent to serve as a new sketch latent that ensembles more visual cues from the given subject, and progressively achieves satisfactory results as the iteration continues. With the aforementioned designs, our contribution can be summarized as follows: 1) We identify that the design choices of the subject image encoder significantly impact identity preservation and propose an improved design over existing solutions. 2) We propose a solution that combines guidance decoupling with iterative bootstrapping to address the guidance balance issue. 3) We extend our approach to personalized image editing and domain-specific personalization tasks. 4) Extensive experiments on both SD2.1 and SDXL demonstrate the state-of-the-art performance and model-agnostic nature of EZIGen.
@@ 2 Related Works
2.2 Tuning-based personalized image generation.
@@ 2.3 Zero-shot personalized image generation
To tackle the aforementioned issues, zero-shot methods were developed for accepting new subjects without retraining. Some[36, 45, 26] are developed for domain-specific tasks, such as human content generation. Recently, for general objects, ELITE[44] pioneeringly projects the image into text space and refines with detailed patch features, BLIP-Diffusion[23] leverages a Q-Former[24] to create multimodal query tokens, and [31, 42, 44] leverage CLIP image encoders to obtain subject representations, BootPIG[33] setups a trainable Reference UNet to extract subject features. Nevertheless, these methods struggle with issues like degraded subject-identity preservation due to sub-optimal feature extraction or poor balancing between the text and subject. Recently, several concurrent studies [46, 40] employ Transformer-based models [22, 1] as a unified framework for various tasks, establishing a new paradigm. However, their performances in personalized generation remain suboptimal.
@@ 3.1 Encoding and injecting subject feature
The identity information of the subject image is often extracted using an image encoder. Current methods use options like CLIP[31], DINO-V2[8]), or a Reference UNet[33] trained based on Stable Diffusion UNet. Leveraging a Stable Diffusion UNet for subject feature extraction is particularly appealing as it naturally aligns with the Stable Diffusion feature space, reducing the challenge of aligning disparate feature spaces. However, this design raises open questions about configuring the feature extractor and integrating subject encoding into the Stable Diffusion UNet. In BootPIG [33], all Reference UNet parameters are fine-tuned, with progressively noised input images fed at each timestep, allowing subject image features to be injected into its Main UNet for generation. However, we identify potential issues: overly-noised features distort subject information, and tuning Stable Diffusion parameters into a Reference UNet can lead to suboptimal subject encoding, leading to degraded subject encoding(Fig. 2 shows some examples). In this paper, we propose a simple yet effective solution: First, we directly take the vanilla Stable Diffusion UNet and conduct one-step denoising on a one-step noised subject image to obtain intermediate subject feature maps from UNet as the subject appearance feature. Second, we again take the same UNet and insert cross-attention adapters in its transformer block to adopt the subject image guidance.
@@ 3.3 Iterative Appearance Transfer
Based on the above analysis and empirical results, we observe that when the sketch latent $x_{t}$ and the subject share similar semantics, such as rough color patterns, the appearance transfer process improves as the adapter establishes a stronger matching within the UNet’s noisy latent space and could thus transfer more subject appearances. To enhance this, we introduce an iterative generation scheme: when we obtain the clean latent $x_{0}$ after the Appearance Transfer Process, we add noise with level $t$ to $x_{0}$ to obtain a new $x_{t}$ as the editable sketch latent for another round of the Appearance Transfer Process, as shown in the bottom part of Fig. 4. This process repeats until the similarity between the newly generated image and the previous image exceeds a threshold, indicating a complete subject transfer as no new visual cues flow in. Integrating the designs above, our model successfully balances subject identity preservation and text adherence, ensuring comprehensive guidance-adhering without compromise. More details in Sec. 8.
@@ 4.1 Implementation details
Benchmark and evaluation. We mainly evaluate EZIGen on two standard benchmarks: DreamBench [35] for personalized image generation, DreamEdit [25] for personalized image editing. For both tasks, we follow Subject Diffusion’s protocol, averaging scores over 6 random runs. Regarding the evaluation metric, we report CLIP-T and DINO scores to assess text adherence and subject identity on DreamBench and we evaluate DINO-sub for foreground subject similarities on DreamEditBench, using SAM [20] for mask extraction. Please refer to Sec. 9 for details. We additionally validate EZIGen on the challenging human content generation benchmark [45] to test the ability to generalize to domain-specific generation tasks, which usually require special design or in-domain training while EZIGen does not. We report ID preservation and prompt consistency following [45, 31]. Finally, we involve human evaluation for the comparisons mentioned above, The percentage scores denote the frequency of the methods’ generation results being chosen as the best within the given sample batch under specific criteria.
Training details and experiment settings. Our training data consists of 200k image pairs constructed from COCO2014 and YoutubeVIS. In COCO2014, we extract 1-4 objects from each target image as subject images, while in YoutubeVIS, we follow [8] to form pairs using frames of the same subject. To preserve SDXL’s high-resolution generation capability, we incorporate an additional 50k high-quality pairs from Subject200k. A visualization of the dataset is provided in Fig. 11. For fair comparisons, the experiments are done with SD2.1-base in most cases unless specified. The image resolution is set to 512×512 for SD2.1-based and 1024×1024 for SDXL. with the AdamW optimizer and a learning rate of $51e-$ 5. During inference, iterations are designed to stop automatically when the newly generated image exhibits a sufficiently high similarity with the image from the previous loop, ensuring efficient convergence and maintaining generation quality.
@@
Ours +SDXL 0.319 0.719 1
  Over.
Flex. ID. Qual.
0.1 0.11 0.07
0.06 0.02 0.02
0.08 0.02 0.03
0.16 0.15 0.11
0.12 0.17 0.17
0.24 0.23 0.29
0.19 0.29 0.31
@@ 4.2 Method comparisons on DreamBench dataset
In Tab. 1 and Fig. 5, we compare our method with recent zero-shot personalized image generation approaches on the DreamBench dataset [35]. Using only a single reference image, our method surpasses MS-Diffusion (0.316 vs. 0.313 CLIP-T; 0.718 vs. 0.671 DINO) and OmniGen (0.316 vs. 0.314; 0.718 vs. 0.684), indicating stronger text-image alignment and subject fidelity. Compared to OmniCtrl, which achieves the highest CLIP-T score among baselines (0.320), our model attains a significantly higher DINO score (0.718 vs. 0.703), reflecting improved identity preservation despite slightly lower text alignment. Furthermore, our SDXL-based variant further pushes the performance to 0.319 (CLIP-T) and 0.719 (DINO), setting a new state-of-the-art across both metrics. These results underscore the effectiveness of our simplified design and the advantages of leveraging the original Stable Diffusion UNet as a feature extractor for zero-shot subject-driven generation.
@@ 4.3 Validation on personalized image editing task
As outlined in Sec. 3.4, our method adapts well to personalized image editing by replacing noise addition with image inversion [30] in the Appearance Transfer Process.
@@
We compare our SD2.1-base model against previous state-of-the-art methods on the DreamEdit benchmark, as shown in Tab. 2 and Fig. 6. Our method significantly outperforms others, primarily due to our advanced subject feature utilization, where the original Stable Diffusion UNet produces high-quality, detailed intermediate appearance features, and the adapter accurately matches them to the sketch latent. Notably, even though MimicBrush [7] leverages a tuned Reference UNet and an extra CLIP image encoder for fine-grained feature extraction, it sometimes struggles to modify areas with large semantic differences from their surroundings. For example, in the last column of Fig. 6, the model has difficulty adjusting the dog face in contrast to the astronaut suit. We attribute this limitation to their masked image modeling training strategy, which aims to recover masked areas based on neighboring tokens.
@@ 4.4 Ablation study
Effectiveness of our subject encoder. To evaluate the quality of the encoded subject feature, we start with a fully noised sketch latent (i.e. random noise) and apply subject guidance at all timesteps, dubbed as “full injection”. As shown in Fig. 7 experiments 1) and 2), compared to the original text-guided generation, the encoded subject representation resembles the subject effectively, achieving a DINO score of 0.762. However, the strong subject features override the weaker sketch latent in early timesteps, leading to severe copy-paste effects and a CLIP-T score of 0.286.
@@
  Subject Encoder Training Strategies DINO Score Dataset Size
@@
Figure 9: Choices of subject encoders, our method is the most efficacious, achieving the highest quality with minimal data requirements. SDU denotes Stable Diffusion UNet.
Figure 10: Evaluating subject ID preservation under different SDU configurations. Left: The impact of different SDU training configurations. The best result is obtained by using the fixed SDU and using a small denoising timestep $t^{\text{sub}}$. Right: Validate the best denoising timestep in SDU for feature extraction.
4.5 Comparing feature extractor with other choices
We evaluate the effectiveness and efficiency of our subject extraction compared to CLIP, DINO, and Reference UNet based methods — Subject Diffusion [31], AnyDoor [8], and BootPIG [33]—as the current best practices for each feature extraction paradigm respectively. For fair comparisons, we ablate the models for AnyDoor and Subject Diffusion (bars 2 and 4) so that subject information is derived solely from the feature extractor. We denote Stable Diffusion UNet, our base extractor, as SDU for simplicity.
Evaluating alternative feature extractor configurations. We examined the impact of tuning SDU parameters and adding varying levels of noise to subject images. Bar chart Fig. 10 shows the experiment settings and their performances, where check marks/cross marks indicate whether the SDU extractor is trained or frozen, the corresponding visualization results are shown in Fig. 10. “ada. time” means that the tuned SDU extractor adaptively shares the same denoising timestep as the main SDU, while arguments starting with “$t^{\text{sub}}$” indicate the level of noise we add to the subject image. From bars (1) and (3), we find that tuning the SDU disrupts the original generative capabilities of the SDU, resulting in a performance drop compared to freezing it. bars (1) and (2) show that fixing the denoising timestep of the SDU to a small $t^{\text{sub}}$ produces better subject identity with clearer feature maps. By fixing the SDU and adjusting $t^{\text{sub}}$, we validated its importance; as shown in bars (5) to (9), DINO scores increase as $t^{\text{sub}}$ decreases. This aligns with the intuition that the SDU encodes clearer features in the final timesteps of the generation process.